OpenStudy (anonymous):

x - 4 / y + 2 = 5 solve for x and show steps please!

OpenStudy (compassionate):

\[\frac{ 2 - 4 }{y + 2 } = 5\] What do you think the first step is?

OpenStudy (anonymous):

why did you change x to 2?

OpenStudy (compassionate):

Oops. I meant x lol

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

well i think the first step is to.. gimme a sec to think, im not to good at these

OpenStudy (anonymous):

to add 4 to 5 i guess?

OpenStudy (anonymous):

@Compassionate ?

OpenStudy (compassionate):

One second.

OpenStudy (anonymous):

oh sorry i thought you were gone for a sec

OpenStudy (compassionate):

cross multiply x - 4 = 5(y + 2) x - 4 = 5y + 10 5y = -x + 14 y = -1/5x + 14/5

OpenStudy (anonymous):

but this is for x, and the answer is showing for y

OpenStudy (compassionate):

-_- Then just do the opposite.

OpenStudy (compassionate):

x - 4 = 5y + 10 Just solve for x.

OpenStudy (anonymous):

i dont know how to do the opposite, i need LOTS of help

OpenStudy (anonymous):

i mean LOTS of help

OpenStudy (anonymous):

so x - 14 = 5y

OpenStudy (anonymous):

@Compassionate ?

OpenStudy (compassionate):

Lol what the heck are you doing xD

OpenStudy (compassionate):

x - 4 = 5y + 10 x = 5y + 14

OpenStudy (anonymous):

idk i cant figure it out!! XD

OpenStudy (anonymous):

ohhhhhhhhhhhhhh

OpenStudy (anonymous):

this was probably the hardest problem ive done in openstudy XD

OpenStudy (anonymous):

but thx for the help